Question 2

Grouper Company purchased a delivery truck for R$34,700 on January 1, 2017. The truck has an expected residual value of R$5,300, and is expected to be driven 100,600 miles over its estimated useful life of 9 years. Actual miles driven were 15,000 in 2017 and 12,200 in 2018.

Calculate depreciable cost per unit. (Round answers to 2 decimal places, e.g. 0.50.)
Depreciable cost per unit R$

per mile
Compute depreciation expense for 2017 and 2018 using the straight-line method, the units-of-activity method, and the double-declining-balance method. (For double-declining-balance round the rate to 2 decimal places, e.g. 22.22%. Round answers to 0 decimal places, e.g. 2,125.)

2017

2018

Straight-line method R$

R$

Units-of-activity method R$

R$

Declining-balance method R$

R$

Assume that Grouper uses the straight-line method.
(1) Prepare the journal entry to record 2017 depreciation.
(2) Show how the truck would be reported in the December 31, 2017, statement of financial position.

(Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

GROUPER COMPANY
Statement of Financial Position

December 31, 2017For the Year Ended December 31, 2017For the Quarter Ended December 31, 2017

R$

Add Less

:

R$
Answer #1
Depreciable cost per unit as per unit-of-activity method =($34,700-$5,300) / 100,600 miles =$0.29 per mile
2017 2018
Straight Line method $3,267 $3,267
($34,700-$5,300)/9 years
Units of activity method $4,350 $3,538
(15,000*$0.29)(12,200*$0.29)
Declining balance method $7,710 $5,997
(Depreciation rate =1/9 years*2 =22.22%)
1
Account and explanation Debit Credit
Depreciation expenses $3,267
Accumulated Depreciation-Equipment $3,267
2 Equipment,Cost $34,700
Less:Accumulated Depreciation $3,267
Equipment,Net $31,433